OpenStudy (anonymous):

Anybody willing to help me with three easy questions please! I will fan and medal and love you forever <3

OpenStudy (anonymous):

If h(x) = 4x − 8 and j(x) = −x, solve h[j(3)] and select the correct answer below. A. −36 B. −20 C. 1 D. 4

OpenStudy (anonymous):

a) what is \(j(3)\)?

OpenStudy (anonymous):

J is -x and 3 is just 3.... wait are u asking me or are you confused? :)

jimthompson5910 (jim_thompson5910):

satellite73 knows the answer s/he is asking you to see if you know it

OpenStudy (anonymous):

\[j(x)=-x\] so \[j(\spadesuit)=-\spadesuit\] and \[j(\xi)=-\xi\] so what is \[j(3)\]?

OpenStudy (anonymous):

Oh ok... Yes so I did 4x-8[-x(3)] but i keep getting 28

OpenStudy (anonymous):

umm... -3?

jimthompson5910 (jim_thompson5910):

j(3) is -3, yes

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

whew b) what is \(h(-3)\)?

jimthompson5910 (jim_thompson5910):

notice how satellite73 replaced the "j(3)" with "-3" going from h[ j(3) ] to h(-3) this only works because j(3) = -3

OpenStudy (anonymous):

4x+24?

jimthompson5910 (jim_thompson5910):

replace every x in h(x) = 4x − 8 with -3

OpenStudy (anonymous):

oh ok so 4(-3) or -12 -8?

OpenStudy (anonymous):

\[h(\diamondsuit)=4\times\diamondsuit-8\]\[h(\Phi)=4\times \Phi -8\]\[h(-3)+?\]

jimthompson5910 (jim_thompson5910):

yeah just compute -12-8

OpenStudy (anonymous):

-20

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

omg tysm!!

OpenStudy (anonymous):

I need help with 2 more... please

jimthompson5910 (jim_thompson5910):

go ahead

OpenStudy (anonymous):

Ray is purchasing a laptop that is on sale for 25% off. He knows the function that represents the sale price of his laptop is c(p) = 0.75p, where p is the original price of the laptop. He also knows he has to pay 8% sale's tax on the laptop. The price of the laptop with tax is f(c) = 1.08c, where c is the sale price of the laptop. Determine the composite function that can be used to calculate the final price of Ray's laptop by solving for f[c(p)]. A. f[c(p)] = 1.83p B. f[c(p)] = 1.83cp C. f[c(p)] = 0.81p D. f[c(p)] = 0.81cp

jimthompson5910 (jim_thompson5910):

\[\Large f(c) = 1.08c\] \[\Large f(\color{red}{c}) = 1.08\color{red}{c}\] \[\Large f(\color{red}{c}) = 1.08(\color{red}{c})\] \[\Large f(\color{red}{c(p)}) = 1.08(\color{red}{c(p)})\] \[\Large f(\color{red}{c(p)}) = ??\]

OpenStudy (anonymous):

um...

jimthompson5910 (jim_thompson5910):

\[\Large f(c) = 1.08c\] \[\Large f(\color{red}{c}) = 1.08\color{red}{c}\] \[\Large f(\color{red}{c}) = 1.08(\color{red}{c})\] \[\Large f(\color{red}{c(p)}) = 1.08(\color{red}{c(p)})\] \[\Large f(\color{red}{c(p)}) = 1.08(\color{red}{0.75p})\] \[\Large f(\color{red}{c(p)}) = ??\] In the second to the last step, I replaced c(p) with 0.75p because c(p) = 0.75p

OpenStudy (anonymous):

0.81

jimthompson5910 (jim_thompson5910):

0.81p

jimthompson5910 (jim_thompson5910):

don't lose the p

jimthompson5910 (jim_thompson5910):

the c or c(p) goes away because it is replaced by 0.75p

OpenStudy (anonymous):

oh ok got it so c right

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

The following table shows the length and width of two rectangles: Length Width Rectangle A 2x + 8 x − 1 Rectangle B 3x + 5 2x − 2 Which expression is the result of the perimeter of rectangle B minus the perimeter of rectangle A? A. 2x − 4 B. 8x + 10 C. 16x + 20 D. 4x − 8

OpenStudy (anonymous):

Thank you! Last question

jimthompson5910 (jim_thompson5910):

hint: Perimeter = 2*(Length + Width)

OpenStudy (anonymous):

ok so, 2*(2x+8)+(x-1) for the first one right?

jimthompson5910 (jim_thompson5910):

2*[ (2x+8)+(x-1) ] actually

OpenStudy (anonymous):

oh ok now what? when i simplify it i get 5x+15

jimthompson5910 (jim_thompson5910):

you should get 2*(3x+7) = ???

OpenStudy (anonymous):

ohh i see! ok so 6x+14?

jimthompson5910 (jim_thompson5910):

yes, that is the perimeter of rect A

jimthompson5910 (jim_thompson5910):

find the perimeter of rect B

OpenStudy (anonymous):

so for rectangle B would be 2*[ (3x+5)+(2x-2) ]

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

I'll be right back

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is the answer 10x+6

jimthompson5910 (jim_thompson5910):

perimeter of rect A = 6x+14 perimeter of rect B = 10x+6

jimthompson5910 (jim_thompson5910):

subtract the two perimeters

OpenStudy (anonymous):

Well the directions say "perimeter of rectangle B minus the perimeter of rectangle A" but when i do that i get 4x+8 which is none of the options

jimthompson5910 (jim_thompson5910):

10x - 6x = 4x ... correct 6 - 14 = 8 ... incorrect

OpenStudy (anonymous):

Now I'm confused

jimthompson5910 (jim_thompson5910):

B - A = (10x+6) - (6x+14) = 10x + 6 - 6x - 14 = ??

OpenStudy (anonymous):

i got 4x+20

OpenStudy (anonymous):

Did you mean to put 10x + 6 - 6x + 14 = ?

jimthompson5910 (jim_thompson5910):

no the negative distributes through

jimthompson5910 (jim_thompson5910):

- (6x+14) = -6x-14

OpenStudy (anonymous):

oh ok so it would be the last one right?

jimthompson5910 (jim_thompson5910):

yeah 4x-8

OpenStudy (anonymous):

Thank you so much!!!!!!!!

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (denisaboichuk):

0